# Unsupervised Evolutionary Cell Type Matching via Entropy-Minimized Optimal Transport

### v1

@@ Abstract
Identifying evolutionary correspondences between cell types across species is a fundamental challenge in comparative genomics and evolutionary biology. Existing approaches often rely on either reference-based matching, which imposes asymmetry by designating one species as the reference, or projection-based matching, which may increase computational complexity and obscure biological interpretability at the cell-type level. Here, we present OT-MESH, an unsupervised computational framework leveraging entropy-regularized optimal transport (OT) to systematically determine cross-species cell type homologies. Our method uniquely integrates the Minimize Entropy of Sinkhorn (MESH) technique to refine the OT plan. It begins by selecting genes with high Signal-to-Noise Ratio (SNR) to capture the most informative features, from which a cost matrix is constructed using cosine distances between cell-type centroids. Importantly, the MESH procedure iteratively refines the cost matrix, leading to a transport plan with significantly enhanced sparsity and interpretability of the resulting correspondence matrices. Applied to retinal bipolar cells (BCs) and retinal ganglion cells (RGCs) from mouse and macaque, OT-MESH accurately recovers known evolutionary relationships and uncovers novel correspondences, one of which was independently validated experimentally. Thus, our framework offers a principled, scalable, symmetric, and interpretable solution for evolutionary cell type mapping, facilitating deeper insights into cellular specialization and conservation across species.
@@ 1 Introduction
Understanding evolutionary relationships between cell types across species is central to elucidating mechanisms underlying cellular specialization and diversity. Recent advances in single-cell RNA sequencing (scRNA-seq) have greatly improved the characterization of cell types based on their gene expression profiles across diverse species wangTracingCelltypeEvolution2021 ; hahnEvolutionNeuronalCell2023a . However, a fundamental challenge persists: given independently identified clusters of transcriptionally similar cells from different species, how can we systematically and reliably determine their evolutionary correspondences?
@@
Existing methods for cross-species cell type alignment broadly fall into two categories shaferCrossSpeciesAnalysisSingleCell2019 : (1) reference-based matching, in which classifiers are trained on labeled data from a reference species to predict cell types in a query species shekharCOMPREHENSIVECLASSIFICATIONRETINAL2016a ; pengMolecularClassificationComparative2019a ; zhangEvolutionaryDevelopmentalSpecialization2024 ; wangMolecularCharacterizationSea2024 ; and (2) projection-based matching, in which individual-cell gene expression profiles from different species are embedded into a shared low-dimensional space using algorithms such as Canonical Correlation Analysis (CCA), Harmony, or other manifold alignment techniques butlerIntegratingSinglecellTranscriptomic2018b ; korsunskyFastSensitiveAccurate2019 ; rosenUniversalCellEmbeddings2024a . Reference-based matching inherently introduces asymmetry by designating one species as the "reference" and the other as the "query," potentially biasing outcomes, especially when species differ significantly in cell type numbers or evolutionary distances. Projection-based matching, though symmetric, typically operates at the level of individual cells, resulting in high computational demands and potentially obscuring direct biological interpretability of cell-type level relationships. Alternatively, cell type correspondences have been inferred through expert-curated morphological, physiological, or marker-gene criteria hahnEvolutionNeuronalCell2023a ; badenVertebrateRetinaWindow2024 . While valuable, these expert-driven methods depend heavily on prior hypotheses and manual validations, limiting scalability and systematic applicability.
To address these limitations, we introduce OT-MESH, an unsupervised computational framework. It combines entropy-regularized optimal transport (OT), a powerful tool for comparing distributions and finding correspondences, with a novel integration of the Minimize Entropy of Sinkhorn (MESH) refinement strategy zhangUnlockingSlotAttention2023 . OT is well-suited for this task due to its inherent ability to handle asymmetric numbers of types and partial correspondences. However, standard entropy-regularized OT often yields diffuse, non-sparse transport plans, hindering direct biological interpretation. Inspired by developments in machine learning and computer vision locatelloObjectCentricLearningSlot2020 ; sanderSinkformersTransformersDoubly2022 , we adapt the MESH technique to iteratively refine the OT cost matrix, promoting sparsity in the cross-species cell type alignment. Our method begins with Signal-to-Noise Ratio (SNR) based gene selection to identify informative features. It then constructs a cost matrix from cosine distances between cell-type centroids in this selected gene space. The MESH procedure is subsequently applied to refine the transport plan, yielding sparse and interpretable correspondence matrices.
@@
We illustrate our approach using established datasets of bipolar cells (BCs) and retinal ganglion cells (RGCs) from mouse and macaque retinas hahnEvolutionNeuronalCell2023a ; pengMolecularClassificationComparative2019a ; tranSinglecellProfilesRetinal2019a . Leveraging existing knowledge for validation, we demonstrate that our method consistently identifies known evolutionary correspondences and discovers novel cell-type homologies, including one correspondence independently validated experimentally wangONtypeDirectionselectiveGanglion2023 . Thus, OT-MESH provides a principled, scalable, and interpretable computational framework, enhancing our ability to map cell type evolution and advance its understanding.
@@ 2.3 OT Applications in Single-Cell Biology
OT methods have found increasing applications in single-cell biology, particularly for aligning cell states across experimental conditions, measurement modalities, or developmental time points demetciSCOTSingleCellMultiOmics2022 ; gossiMatchingSingleCells2023 ; penaConstructingCelltypeTaxonomy2025 . A key advantage of OT approaches lies in their natural ability to handle partial or incomplete correspondences. Recent studies have leveraged OT to construct hierarchical cell-type taxonomies, effectively identifying cell types uniquely represented in specific datasets penaConstructingCelltypeTaxonomy2025 . Nonetheless, OT-based methods have yet to be comprehensively developed and validated specifically for cross-species evolutionary cell type comparisons. In this study, we address this critical gap by adapting and extending OT methodologies, particularly through the MESH procedure, to the problem of cross-species cell type matching.
@@ 3.3 Cost Matrix Construction
We define a scale-invariant cost matrix $\mathbf{C}\in\mathbb{R}^{a\times b}$ based on the cosine distance between cell-type centroids:
@@ 3.5 Algorithm Summary
Figure 1: Illustration of the proposed method. Cell type centroids are computed using SNR-selected genes, and a cost matrix is constructed using cosine distance between these centroids for the first and second species. The MESH procedure iteratively refines this cost matrix to produce the final correspondence matrix that reveals evolutionarily related cell types with sparse, interpretable mappings that quantify the strength of relationships between cell types across species.
@@ 4.1 Datasets
We evaluated our method using three established retinal cell datasets:
@@ 4.3.1 Evaluation Metrics
  • Sparseness Score: Proportion of negligible entries (<0.001) in the correspondence matrix, indicating interpretability through sparsity.
@@ 4.3.2 Parameter Selection
We conducted systematic parameter optimization through grid search over the following parameter ranges:
  • Regularization parameter $\alpha$: [0.01, 0.1, 1.0, 10.0]
  • MESH learning rate $\lambda$: [0.1, 0.5, 1.0, 5.0]
  • Number of MESH iterations $T$: [1.0, 2.0, 4.0, 8.0]
For each parameter combination, we computed the transport loss (Eq. 2) and sparseness score. Because our primary aim was sparse and interpretable correspondences, we restricted our selection to parameter sets with sparseness scores above 0.5, choosing the set that minimized transport loss. Optimal parameters were independently determined for each experiment, detailed in SupplementaryFigures 1–3.
@@ 5 Results
5.1 Method Validation and Comparison
@@
We first validated our OT-MESH method against known cell type correspondences between macaque peripheral and foveal bipolar cells (BCs), comparing its performance to three baseline methods: XGBoost, Harmony+1NN, and standard entropy-regularized OT without MESH. For XGBoost, the classifier was trained using macaque foveal BCs as the reference pengMolecularClassificationComparative2019a . For Harmony+1NN, we performed embedding based integration followed by cell type label transfer through nearest neighbor assignment in the harmonized PCA space korsunskyFastSensitiveAccurate2019 . Standard OT was performed using the same parameters as OT-MESH but without iterative entropy minimization.
@@
Figure 2 illustrates the resulting correspondence matrices. The XGBoost approach (Figure 2A) yielded a predominantly diagonal structure with notable off-diagonal noise. Harmony+1NN (Figure 2B) improved upon XGBoost by producing cleaner correspondences, though some off-diagonal elements persisted. Standard OT without MESH refinement (Figure 2C) produced a highly diffuse correspondence matrix, making interpretation nearly impossible despite its computational speed. In contrast, OT-MESH (Figure 2D) resulted in a clean diagonal structure with high-confidence values, clearly recovering the true 1:1 correspondences.
Quantitative metrics (Table 1) reinforce these findings. OT-MESH achieved the best overall performance: highest sparseness (0.8678), lowest entropy (2.4597), and highest ARI (0.9779). Harmony+1NN was second in sparseness (0.8347) and ARI (0.9258) but was significantly slower (14.5306s vs. 0.7626s for OT-MESH). XGBoost had lower ARI (0.8871) and was slower than OT-MESH (2.3062s). Standard OT was fastest (0.0329s) but its extremely low sparseness (0), poor ARI (0.0240), and diffuse nature made it unsuitable for meaningful biological interpretation. These results demonstrate that OT-MESH provides a compelling balance of accuracy, interpretability, and computational efficiency for cross-species cell type correspondence.
@@
Figure 2: Comparison of correspondence matrices for macaque peripheral to foveal BC type mapping. A) XGBoost shows diagonal structure but with some off-diagonal noise. B) Harmony+1NN improves alignment with cleaner structure. C) Regular OT produces a diffuse, uninterpretable mapping. D) OT-MESH yields the clearest diagonal structure with high-confidence 1:1 matchings, consistent with ground truth. Correspondence matrices are normalized so that the sum of all their entries is one.
Table 1: Performance comparison of cross-species cell type matching methods on macaque foveal vs. peripheral BC types. Best values for each metric are in bold.
Method Sparseness Score Entropy ARI Running Time (s)
XGBoost 0.8017 2.6738 0.8871 2.3062
Harmony+1NN 0.8347 2.5949 0.9258 14.5306
OT without MESH 0.0000 4.6934 0.0240 0.0329
OT-MESH 0.8678 2.4597 0.9779 0.7626
5.2 Cross-Species Correspondence of Bipolar Cell (BC) Types

### v2

@@ Abstract
Identifying evolutionary correspondences between cell types across species is a fundamental challenge in comparative genomics and evolutionary biology. Existing approaches often rely on either reference-based matching, which imposes asymmetry by designating one species as the reference, or projection-based matching, which may increase computational complexity and obscure biological interpretability at the cell-type level. Here, we present OT-MESH, an unsupervised computational framework leveraging entropy-regularized optimal transport (OT) to systematically determine cross-species cell type homologies. Our method uniquely integrates the Minimize Entropy of Sinkhorn (MESH) technique to refine the OT plan, transforming diffuse transport matrices into sparse, interpretable correspondences. Through systematic evaluation on synthetic datasets, we demonstrate that OT-MESH achieves near-optimal matching accuracy with computational efficiency, while maintaining remarkable robustness to noise. Compared to other OT-based methods like RefCM, OT-MESH provides speedup while achieving comparable accuracy. Applied to retinal bipolar cells (BCs) and retinal ganglion cells (RGCs) from mouse and macaque, OT-MESH accurately recovers known evolutionary relationships and uncovers novel correspondences, one of which was independently validated experimentally. Thus, our framework offers a principled, scalable, and interpretable solution for evolutionary cell type mapping, facilitating deeper insights into cellular specialization and conservation across species.
@@ 1 Introduction
Understanding evolutionary relationships between cell types across species is central to elucidating mechanisms underlying cellular specialization and diversity. Recent advances in single-cell RNA sequencing (scRNA-seq) have enabled unprecedented characterization of cell types based on their gene expression profiles across diverse species wangTracingCelltypeEvolution2021 ; hahnEvolutionNeuronalCell2023a , with atlas-scale datasets now containing hundreds to thousands of distinct cell types. However, a fundamental challenge persists: given independently identified clusters of transcriptionally similar cells from different species, how can we systematically, reliably, and efficiently determine their evolutionary correspondences?
@@
Existing methods for cross-species cell type alignment broadly fall into two categories shaferCrossSpeciesAnalysisSingleCell2019 : (1) reference-based matching, in which classifiers are trained on labeled data from a reference species to predict cell types in a query species shekharCOMPREHENSIVECLASSIFICATIONRETINAL2016a ; pengMolecularClassificationComparative2019a ; zhangEvolutionaryDevelopmentalSpecialization2024 ; wangMolecularCharacterizationSea2024 ; and (2) projection-based matching, in which individual-cell gene expression profiles from different species are embedded into a shared low-dimensional space using algorithms such as Canonical Correlation Analysis (CCA), Harmony, or other manifold alignment techniques butlerIntegratingSinglecellTranscriptomic2018b ; korsunskyFastSensitiveAccurate2019 ; rosenUniversalCellEmbeddings2024a .
Reference-based matching inherently introduces asymmetry by designating one species as the "reference" and the other as the "query," potentially biasing outcomes, especially when species differ significantly in cell type numbers or evolutionary distances. Projection-based matching, though symmetric, typically operates at the level of individual cells, resulting in high computational demands and potentially obscuring direct biological interpretability of cell-type level relationships. Recent optimal transport (OT) based methods like RefCM galantiAutomatedCellType2024 offer improved theoretical foundations but rely on computationally expensive exact earth mover’s distance (EMD) calculations, limiting their scalability for large atlases.
To address these limitations, we introduce OT-MESH, an unsupervised computational framework that combines entropy-regularized OT with the Minimize Entropy of Sinkhorn (MESH) refinement strategy zhangUnlockingSlotAttention2023 . The key insight is that while standard entropy-regularized OT is computationally efficient but produces diffuse, non-sparse transport plans unsuitable for biological interpretation. The MESH technique iteratively refines the cost matrix to promote sparsity in the cross-species cell type alignment while maintaining the computational advantages of entropy regularization. This approach enables OT-MESH to achieve the interpretability while maintaining the speed necessary for large-scale comparative studies.
@@
We illustrate our approach through comprehensive experiments on both synthetic and biological datasets. Using synthetic data with known ground truth correspondences, we demonstrate that OT-MESH maintains high accuracy with sub-second runtime for up to 500 cell types—54× faster than RefCM while achieving comparable performance. The method shows remarkable robustness, maintaining high matching accuracy even at 8× noise levels, substantially outperforming projection-based and supervised approaches. Applied to established datasets of bipolar cells (BCs) and retinal ganglion cells (RGCs) from mouse and macaque retinas hahnEvolutionNeuronalCell2023a ; pengMolecularClassificationComparative2019a ; tranSinglecellProfilesRetinal2019a , OT-MESH consistently identifies known evolutionary correspondences and discovers novel cell-type homologies, including one correspondence independently validated experimentally wangONtypeDirectionselectiveGanglion2023 , demonstrating the method’s biological relevance. Thus, OT-MESH provides a principled, scalable, and interpretable computational framework, enhancing our ability to map cell type evolution and advance its understanding.
@@ 2.3 OT Applications in Single-Cell Biology
OT methods have found increasing applications in single-cell biology, particularly for aligning cell states across experimental conditions, measurement modalities, or developmental time points demetciSCOTSingleCellMultiOmics2022 ; gossiMatchingSingleCells2023 ; galantiAutomatedCellType2024 ; penaConstructingCelltypeTaxonomy2025 . A key advantage of OT approaches lies in their natural ability to handle partial or incomplete correspondences, and recent studies have leveraged OT to construct hierarchical cell-type taxonomies or annotations galantiAutomatedCellType2024 ; penaConstructingCelltypeTaxonomy2025 . Specifically, Galanti et al. galantiAutomatedCellType2024 introduced reference cluster mapping (RefCM), which combines optimal transport with integer programming to enforce sparse mappings between reference and query datasets. This approach addresses the over-diffusion problem in standard OT by solving a secondary discrete optimization problem, allowing for complex relationships including many-to-one and one-to-many mappings between cell types. While RefCM and other OT-based methods have shown promise for within-species cell type annotation, they have yet to be comprehensively validated specifically for cross-species evolutionary cell type comparisons. In this study, we address this gap by adapting and extending OT methodologies, particularly through the MESH procedure, to the problem of cross-species cell type matching.
@@ 3.3 Cost Matrix Construction
We define a cost matrix $\mathbf{C}\in\mathbb{R}^{a\times b}$ to quantify the dissimilarity between cell types across species. This cost is based on the cosine distance between the high-dimensional gene expression centroids:
@@
Our choice of cosine distance is deliberate. In cross-species comparisons, homologous cell types may preserve the relative pattern of their gene expression profiles but exhibit significant differences in the absolute magnitude of expression due to technical variability (e.g., sequencing depth) or biological divergence. Cosine distance is critically insensitive to these magnitude shifts, as it only measures the angle between two vectors. In contrast, a metric like Euclidean distance would be confounded by such scaling effects, potentially assigning a high dissimilarity cost to genuinely related cell types. Thus, cosine distance provides a robust measure of profile similarity.
@@ 3.5 Algorithm Summary
Figure 1: Overview of the OT-MESH framework for cross-species cell type matching. Starting from cell type centroids computed using SNR-selected genes, a cost matrix is constructed based on cosine distances between centroids from two species. Standard entropy-regularized OT yields an initial correspondence matrix $\mathbf{W}$ that is typically diffuse and difficult to interpret biologically. The MESH procedure iteratively refines the cost matrix through entropy minimization, transforming the diffuse transport plan into a sparse, interpretable correspondence matrix $\mathbf{W}^{*}$ that clearly identifies evolutionarily related cell types between species.
@@ 4.1 Datasets
4.1.1 Synthetic Data for Scalability and Robustness Analysis
To systematically evaluate computational scalability and noise robustness, we generated synthetic matched datasets simulating cross-species correspondence scenarios. These datasets consisted of paired "species" with known 1:1 correspondences, where expression profiles were generated using housekeeping genes (expressed across all types) and cell-type-specific marker genes, with Poisson-sampled counts and controlled noise levels. For scalability testing, we varied the number of cell types (50, 100, 200, 500 types), while for robustness analysis, we fixed the dataset at 50 cell types and varied noise levels (1.0, 2.0, 4.0 ,8.0). All synthetic experiments used identity matrices as ground truth, enabling precise accuracy quantification through ARI. Detailed information about the data generation process can be found in Appendix 9.1.
4.1.2 Retinal Cell Datasets
We also evaluated our method using three established retinal cell datasets:
@@ 4.3.1 Evaluation Metrics
  • Sparsity Score: Proportion of negligible entries (<0.001) in the correspondence matrix, indicating interpretability through sparsity.
@@ 4.3.2 Parameter Selection
While OT-MESH involves three main hyperparameters: regularization strength ($\alpha$), MESH learning rate ($\lambda$), and number of MESH iterations ($T$), we developed a parameter selection strategy based on the observation that $\lambda$ and $T$ primarily control the sparsity of the resulting transport plan. Our selection heuristic uses an elbow method to identify parameter combinations that achieve sufficient sparsity without over-regularization. Specifically, for each $\alpha$ value, we visualize the relationship between the entropy and different values of $\lambda$ and $T$ to identify the elbow point where the plot shows a sharp bend, indicating diminishing effects of reducing entropy with adjusting parameters (SupplementaryFigure 1A,B).
From this set of high-sparsity candidates (one for each $\alpha$), we select the final parameter triplet (including $\alpha$) that minimizes the transport cost, thus balancing sparsity with fidelity to the cost matrix (SupplementaryFigure 1C). This approach provides data-driven method for parameter selection that ensures reproducibility while allowing flexibility to adapt to different dataset characteristics.
To apply this heuristic efficiently across datasets of varying scales, we utilize adaptive search ranges. The parameter ranges are defined as follows:
  • For smaller-scale problems (< 100 types per species, e.g., retinal cell type comparisons):
    • Regularization parameter $\alpha$: [0.1, 0.5, 1.0, 5.0, 10.0]
    • MESH learning rate $\lambda$: [0.1, 0.5, 1.0, 5.0, 10.0]
    • Number of MESH iterations $T$: [2.0, 4.0, 6.0, 8.0, 10.0]
  • For larger-scale problems ($\geq$ 100 types per species, e.g., scalability analysis): the search ranges are extended to [0.5, 1.0, 5.0, 10.0, 50.0] for $\lambda$ and to [8.0, 10.0, 12.0, 14.0, 16.0] for $T$, as for larger and more complex transport matrices, a more aggressive optimization schedule is required to effectively drive the solution towards a sparse and interpretable state.
@@ 4.3.3 Comparative Baselines
  • RefCM: An OT-based method galantiAutomatedCellType2024 that maps query cell type clusters to reference clusters through controlled complexity parameters: max_merges (maximum query clusters mapping to one reference cluster), max_splits (maximum reference clusters per query cluster), and discovery_threshold (cost threshold above which clusters are considered novel types rather than mapped). We evaluated two configurations to assess performance under different mapping constraints: RefCM-Strict (max_merges/splits=1, discovery_threshold=0.0) enforcing strict 1-to-1 mappings comparable to our expected ground truth, and RefCM-Flexible (max_merges/splits = 10% of cell types, discovery_threshold=0.5) permitting complex many-to-many correspondences to test robustness when the true mapping structure is unknown.
@@ 5 Results
5.1 Application to Synthetic Datasets
5.1.1 Computational Scalability Analysis
To evaluate the performance of OT-MESH, especially its scalability, and compare it against baseline methods, we conducted systematic benchmarks using synthetic datasets with varying numbers of cell types (50, 100, 200, and 500 types). Each dataset maintained 50 cells per type with known 1:1 correspondences, allowing precise quantification of matching accuracy.
Figure 2 presents the scaling behavior across three key metrics. OT-MESH demonstrated favorable computational scaling (Figure 2A), maintaining sub-second runtime across all tested scales (0.013s for 50 types to 0.113s for 500 types). In contrast, XGBoost showed steeper scaling: its runtime increased from 1.052s for 50 types to 60.53s for 500 types due to the growing complexity of multi-class classification. Harmony+1NN exhibited similar challenges, requiring 3.216s for 50 types and 18.266s for 500 types due to the computational cost of embedding and nearest-neighbor search. On average, both RefCM variants showed the highest computational costs among all methods, requiring approximately 160s for the largest dataset due to the complexity of computing optimal transport over earth mover’s distances (EMDs).
For matching accuracy (Figure 2B), OT-MESH maintained consistently high performance (ARI > 0.69) across all scales. RefCM-Strict achieved perfect accuracy (ARI = 1.0) as expected, given that it enforces 1-to-1 mappings matching our ground truth structure, though at substantial computational cost. When this constraint was relaxed in RefCM-Flexible, performance degraded significantly, with ARI declining from 0.4 to 0.2 as problem size increased. Both XGBoost and Harmony+1NN underperformed relative to OT-MESH, while standard entropy-regularized OT without MESH completely failed (ARI close to 0) across all scales, confirming the necessity of entropy minimization in our approach.
The entropy analysis (Figure 2C) revealed that OT-MESH solutions maintained low entropy (4.6-6.8) across all scales, indicating concentrated, interpretable correspondence matrices. Standard OT without MESH produced highly diffuse solutions with the highest entropy values, increasing from 7.8 to 12.4 as the number of types grew. RefCM-Strict achieved the lowest entropy values (3.9-6.2), consistent with its enforced 1-to-1 constraint, while other methods showed intermediate values.
Figure 2: Scalability analysis across varying numbers of cell types. A) Runtime scaling of different methods. B) Matching accuracy (ARI) over cell type number of different methods. C) Solution entropy scaling of different methods. Different methods are indicated in the legend of panel C. Error bars represent standard deviation across three independent runs.
5.1.2 Robustness to Noise
We evaluated method robustness by systematically varying noise levels (1.0, 2.0, 4.0, and 8.0) on datasets with 50 cell types. Performance degradation patterns across methods are presented in SupplementaryFigure 2.
For matching accuracy (SupplementaryFigure 2A), RefCM-Strict maintained perfect performance (ARI = 1.0) across all noise levels, benefiting from its enforced correspondence structure. OT-MESH demonstrated remarkable robustness with only gradual degradation, maintaining ARI > 0.93 even at the highest noise level. In contrast, XGBoost exhibited steep performance decline, with ARI dropping from 0.67 to 0.05 as noise increased. Harmony+1NN and RefCM-Flexible displayed intermediate robustness, while standard OT without MESH remained ineffective (ARI close to 0) regardless of noise level.
The entropy analysis under varying noise conditions (SupplementaryFigure 2B) revealed distinct response patterns. OT-MESH and RefCM-Strict maintained relatively stable entropy values across all noise levels, indicating that their sparse solutions remained robust to data quality degradation. Conversely, RefCM-Flexible, XGBoost, and Harmony+1NN exhibited increasing entropy with higher noise levels, suggesting their correspondence matrices became progressively more diffuse and uncertain as data quality decreased. Standard OT maintained consistently high entropy across all conditions, reflecting its inherently diffuse nature.
These noise robustness experiments, combined with our scalability analysis, demonstrate that OT-MESH achieves an optimal balance of desirable properties: near-optimal accuracy without requiring prior knowledge of correspondence structure, sub-second runtime even for large-scale problems, and robust performance under challenging noise conditions. While RefCM-Strict guarantees perfect matching when the true correspondence structure is known a priori, OT-MESH provides a practical and efficient solution for real-world scenarios where such knowledge is unavailable.
5.2 Application to Retinal Cell Datasets
5.2.1 Method Validation and Comparison
@@
Applying to the retinal cell datasets, we first validated OT-MESH against known cell type correspondences between macaque peripheral and foveal bipolar cells (BCs), comparing its performance to baseline methods. For XGBoost, the classifier was trained using macaque foveal BCs as the reference pengMolecularClassificationComparative2019a . For Harmony+1NN, we performed embedding-based integration followed by cell type label transfer through nearest neighbor assignment in the harmonized PCA space korsunskyFastSensitiveAccurate2019 . RefCM-Strict was configured to enforce 1-to-1 mappings galantiAutomatedCellType2024 . Standard OT was performed using the same parameters as OT-MESH but without iterative entropy minimization.
@@
Figure 3 illustrates the resulting correspondence matrices. XGBoost (Figure 3A) yielded a predominantly diagonal structure with notable off-diagonal noise. Harmony+1NN (Figure 3B) improved upon XGBoost by producing cleaner correspondences, though some off-diagonal elements persisted. Standard OT without MESH refinement (Figure 3C) produced a highly diffuse correspondence matrix, making interpretation impossible despite its computational speed. In contrast, OT-MESH (Figure 3D) achieved a clean diagonal structure with high-confidence values, effectively recovering the true 1:1 correspondences without requiring any prior constraints. RefCM-Strict (Figure 3E) also achieved perfect diagonal alignment through its enforced 1-to-1 mappings.
Quantitative metrics (SupplementaryTable 1) reveal a striking finding: OT-MESH achieved identical perfect performance to RefCM-Strict (sparsity: 0.9091, entropy: 2.3979, ARI: 1.0000) while being 54× faster (0.6375s vs. 34.4387s). This demonstrates that the MESH procedure effectively discovers the optimal sparse correspondence structure through entropy minimization alone, without requiring explicit structural constraints. Harmony+1NN showed good performance (ARI: 0.9258) but required 23× longer runtime than OT-MESH. XGBoost exhibited lower accuracy (ARI: 0.8871) and was 3.6× slower. While standard OT was fastest (0.0296s), its extremely low sparsity, negligible ARI, and diffuse nature rendered it unsuitable for biological interpretation. These results establish that OT-MESH uniquely combines the accuracy of constrained methods with superior computational efficiency, providing an effective solution for cross-species cell type correspondence.
@@
Figure 3: Comparison of correspondence matrices for macaque peripheral to foveal BC type mapping. A) XGBoost shows diagonal structure with notable off-diagonal noise. B) Harmony+1NN improves alignment with cleaner structure. C) Standard OT produces a diffuse, uninterpretable mapping. D) OT-MESH yields perfect diagonal structure with high-confidence 1:1 matchings without requiring prior constraints. E) RefCM-Strict achieves clean diagonal through enforced 1-to-1 constraints. Correspondence matrices are normalized so that the sum of all their entries is one.
5.2.2 Cross-Species Correspondence of Bipolar Cell (BC) Types